# On the kernels of the pro-$p$ outer Galois representations associated to once-punctured CM elliptic curves

### v2

@@ 1. Introduction
  2. (2)
    The first and second assumptions of Theorem A allow us to determine the $\Lambda$-module structure of the Galois group $\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}$, where
    $\Lambda\coloneqq\mathbb{Z}_{p}[[\mathrm{Gal}(K(p^{\infty})/K(p))]]$
    is the completed group ring over $\mathbb{Z}_{p}$ which is (non-canonically) isomorphic to the two-variable power series ring over $\mathbb{Z}_{p}$. More precisely, under these assumptions, we have
    $\mathrm{Gal}(\Omega/K(p^{\infty}))^{\mathrm{ab}}\cong\Lambda^{[K(p):K]-1}\oplus\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1)),$
    where $\mathrm{ann}_{\Lambda}(\mathbb{Z}_{p}(1))\subset\Lambda$ is the annihilator ideal of the Tate twist $\mathbb{Z}_{p}(1)$, cf. Remark 4.3.
  3. (3)
    We do not know whether there are infinitely many primes satisfying the first assumption of Theorem A, because this assumption for $p$ implies that $p$ is regular. We also do not know whether the second assumption is satisfied by infinitely many primes. For example, there are 39175 prime numbers which are less than $10^{6}$ and split in $\mathbb{Q}(\sqrt{-1})$. Among them, we have checked that 13705 primes satisfy the second assumption.
Example 1.1.
We have checked whether a small prime $p$ satisfies the first two assumptions of Theorem A, by using SageMath [The24]. For example, when $p=5$, both assumptions are satisfied if and only if $K=\mathbb{Q}(\sqrt{-1})$. Similarly, when $p=7$, both conditions are satisfied only for $K=\mathbb{Q}(\sqrt{-3})$ or $\mathbb{Q}(\sqrt{-19})$. When $p=13$ or $17$, we also checked that both assumptions are satisfied for $K=\mathbb{Q}(\sqrt{-1})$, under the Generalized Riemann Hypothesis (GRH).
Our strategy to prove Theorem A is to generalize Sharifi’s technique developed in [Sha02]. That is, we construct $\{\sigma_{\boldsymbol{m}}\}_{\boldsymbol{m}\in I}$ satisfying the assumption of Theorem A in a group-theoretic way such that it generates $\mathrm{Gal}(\Omega/K(p^{\infty}))$, and we then use Conjecture 2.8 to conclude that the surjection $\mathrm{Gal}(\Omega/K(p^{\infty}))\to\mathrm{Gal}(\Omega^{\ast}/K(p^{\infty}))$ is an isomorphism.
To construct such elements, we need to introduce certain two-variable filtrations on the pro-$p$ geometric fundamental group $\Pi_{1,1}$ of $X$, on subgroups of the pro-$p$ mapping class group $\Gamma_{1,1}$ of type $(1,1)$ and on the absolute Galois group $G_{K}$, and establish their fundamental properties in Section 3.
As was mentioned in Remark 1.3 (1), the Galois group $\mathrm{Gal}(\Omega/K(p))$ has a nontrivial relation for every prime $p\geq 5$ which splits in $K$. This is one of the big differences from the previous result on the thrice-punctured projective line (cf. Remark 1.2), making it difficult to extend Sharifi’s approach to the case of once-punctured elliptic curves. Fortunately, however, the Galois group $\mathrm{Gal}(\Omega/K(p^{\infty}))$ is a free pro-$p$ group (under the assumption of Theorem A), cf. Remark 4.3 (2). Therefore, once we obtain an explicit set of generators of $\mathrm{Gal}(\Omega/K(p^{\infty}))$, the rest of the proof follows in the same way as the previous result.
This paper proceeds as follows: In Section 2, we explain previous studies on the pro-$p$ outer Galois representations associated to once-punctured elliptic curves, especially the construction of a certain power series and Nakamura’s explicit formula for the power series. We then explain the properties of certain Kummer characters associated to the power series, propose an analogue of the Deligne-Ihara conjecture and state the main result in a precise manner. In Section 3, we introduce a certain two-variable version of the descending central series on various profinite groups and establish their fundamental properties. Section 4 is devoted to the proof of the main result.
Acknowledgement. This paper is a revised version of a part of the author’s doctoral thesis submitted to Kyoto University. The author sincerely thanks his advisor Akio Tamagawa for helpful comments, and Benjamin Collas for carefully reading drafts of the thesis. He also sincerely appreciates the referees for their constructive comments, which have greatly contributed to improving the content of this paper.
Notation
Indexes. For a pair $\boldsymbol{m}=(m_{1},m_{2})$ in $\mathbb{Z}^{2}$, we write $\lvert\boldsymbol{m}\rvert\coloneqq m_{1}+m_{2}$. Moreover, for an integer $w$, we write $\boldsymbol{m}\equiv 0\bmod w$ if $\boldsymbol{m}\in(w\mathbb{Z})^{2}$. For any two pairs $\boldsymbol{m}=(m_{1},m_{2})$ and $\boldsymbol{n}=(n_{1},n_{2})$, we write $\boldsymbol{m}\geq\boldsymbol{n}$ if $m_{i}\geq n_{i}$ holds for every $i=1,2$. Moreover, we write $\boldsymbol{m}>\boldsymbol{n}$ if $\boldsymbol{m}\geq\boldsymbol{n}$ and $\lvert\boldsymbol{m}\rvert>\lvert\boldsymbol{n}\rvert$. Finally, we often denote $(1,1)$ by $\boldsymbol{1}$.
Profinite Groups. For a profinite group $G$ and a subset $S$ of $G$, let $\langle S\rangle$ denote the closed subgroup of $G$ topologically generated by $S$. We say that $S$ generates $G$ if $G=\langle S\rangle$. Moreover, we say that $S$ strongly111 Note that the term “strong” is specific to this paper. In the book of Ribes-Zalesskii [RZ10, 2.4], the same notion is described as “$S$ is a generating set of $G$ converging to $1$”. generate $G$ if $S$ generates $G$ and $S$ converges to $1$, i.e. every open subgroup of $G$ contains all but a finite number of elements of $S$.
Let $\langle S\rangle_{{\rm normal}}$ denote the minimal normal closed subgroup of $G$ containing $S$. We call $\langle S\rangle_{{\rm normal}}$ the normal closure of $S$. We say that $S$ normally generates $G$ if $\langle S\rangle_{{\rm normal}}$ coincides with $G$.
The descending central series $\{G(m)\}_{m\geq 1}$ of $G$ is defined by
where, for every pair of subgroups $H,K\subset G$, the symbol $[H,K]$ denotes the topological closure of the commutator subgroup of $H$ and $K$. Note that the quotient $G^{\rm{ab}}=G/G(2)$ is the maximal abelian quotient of $G$, which we also denote by $G^{\rm{ab}}$. Moreover, for a prime number $p$, we denote the maximal pro-$p$ quotient of $G$ by $G^{(p)}$.
We denote the (continuous) automorphism group of $G$ by $\mathrm{Aut}(G)$, the inner automorphism group by $\mathrm{Inn}(G)$ and the outer automorphism group of $G$ by $\mathrm{Out}(G)\coloneqq\mathrm{Aut}(G)/\mathrm{Inn}(G)$. For $g\in G$, the associated inner automorphism is denoted by $\mathrm{inn}(g)$.
In this paper, we frequently use free pro-$p$ groups of countably infinite rank. We refer to Ribes-Zalesskii [RZ10, Lemma 3.3.4] for the characterization of free pro-$p$ groups on sets converging to $1$.
Number Fields. Throughout this paper, we work with a fixed algebraic closure $\bar{\mathbb{Q}}$ of $\mathbb{Q}$ and a fixed embedding from $\bar{\mathbb{Q}}$ into $\mathbb{C}$. Every number field is considered to be a subfield of $\bar{\mathbb{Q}}$. For an integer $m\geq 1$, we denote the group of $m$-th roots of unity in $\bar{\mathbb{Q}}$ by $\mu_{m}$.
For a subfield $F$ of $\bar{\mathbb{Q}}$, we denote the absolute Galois group $\mathrm{Gal}(\bar{\mathbb{Q}}/F)$ of $F$ by $G_{F}$ and the ring of integers in $F$ by $O_{F}$. Moreover, for a place $v$ of $F$, we denote the $v$-adic completion of $F$ by $F_{v}$.
For an imaginary quadratic field $K$ and a nonzero integral ideal $\mathfrak{m}$ of $K$, we denote the ray class field of $K$ modulo $\mathfrak{m}$ by $K(\mathfrak{m})$ and write $K(\mathfrak{m}^{\infty})\coloneqq\cup_{n\geq 1}K(\mathfrak{m}^{n})$. If $\alpha\in O_{K}$ generates the ideal $\mathfrak{m}$, we sometimes denote $K(\mathfrak{m})$ and $K(\mathfrak{m}^{\infty})$ by $K(\alpha)$ and $K(\alpha^{\infty})$, respectively.
Elliptic Curves with Complex Multiplication (cf. Silverman [Sil94, Chapter II]). Let $K$ be an imaginary quadratic field of class number one, and $(E,O)$ an elliptic curve over $K$ with its origin $O\in E(K)$ having complex multiplication by $O_{K}$.
For an ideal $\mathfrak{m}$ of $O_{K}$, we denote the $\mathfrak{m}$-torsion subgroup scheme of $E$ by $E[\mathfrak{m}]$. The $G_{K}$-action on $E[\mathfrak{m}](\bar{\mathbb{Q}})$ determines an injective homomorphism
$\mathrm{Gal}(K(E[\mathfrak{m}])/K)\hookrightarrow\mathrm{Aut}(E[\mathfrak{m}](\bar{\mathbb{Q}}))\cong(O_{K}/\mathfrak{m})^{\times}.$
Moreover, this homomorphism induces an isomorphism
$\mathrm{Gal}(K(\mathfrak{m})/K)\xrightarrow{\sim}(O_{K}/\mathfrak{m})^{\times}/\mathrm{im}(O_{K}^{\times}),$
which does not depend on the choice of $E$.
[…]
@@
$\chi_{1}\colon G_{K}\to\mathrm{Aut}(T_{\mathfrak{p}}(E))\cong\mathbb{Z}_{p}^{\times}\quad\text{and}\quad\chi_{2}\colon G_{K}\to\mathrm{Aut}(T_{\bar{\mathfrak{p}}}(E))\cong\mathbb{Z}_{p}^{\times}.$
@@
Note that $\chi_{1}\chi_{2}=\chi_{\mathrm{cyc}}$, the $p$-adic cyclotomic character. For $\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}^{2}$, we define
$\chi^{\boldsymbol{m}}\coloneqq\chi_{1}^{m_{1}}\chi_{2}^{m_{2}}\colon G_{K}\to\mathbb{Z}_{p}^{\times},$
which factors through $\mathrm{Gal}(K(p^{\infty})/K)$ if $m_{1}\equiv m_{2}\bmod\lvert O_{K}^{\times}\rvert$. For a $\mathbb{Z}_{p}$-module $M$ on which the Galois group $G_{K}$ acts continuously, we denote the $\chi^{\boldsymbol{m}}$-twist of $M$ by $M(\boldsymbol{m})$. Using this notation, we have $T_{\mathfrak{p}}(E)\cong\mathbb{Z}_{p}(1,0)$, $T_{\bar{\mathfrak{p}}}(E)\cong\mathbb{Z}_{p}(0,1)$ and $\mathbb{Z}_{p}(m,m)$ is simply the $m$-th Tate twist $\mathbb{Z}_{p}(m)$.
@@ 2. Preliminaries
In this section, we prepare backgrounds which are necessary to state the main result of this paper (Theorem 2.12). In the following, we explain:
  • construction of a certain basis $\{x_{1},x_{2}\}$ of the pro-$p$ geometric fundamental group of once-punctured elliptic curve $X$, cf. Lemma 2.1,
  • construction of an elliptic analogue of Ihara’s universal power series for Jacobi sums [Iha86], and Nakamura’s explicit formula of that power series [Nak95], cf. Section 2.1,
  • conditional nonvanishing and surjectivity of certain Kummer characters associated to the power series [Ish25], cf. Theorem 2.5 and Theorem 2.6,
  • formulation of an analogue of the Deligne-Ihara conjecture, and state the main result, cf. Section 2.2.
In the following, let $K$ be an imaginary quadratic field of class number one, and $p\geq 5$ a prime which splits into two distinct primes in $O_{K}$ as $(p)=\mathfrak{p}\bar{\mathfrak{p}}$. Let $(E,O)$ be an elliptic curve defined over $K$ having complex multiplication by $O_{K}$ with a Weierstrass form of $y^{2}=4x^{3}-g_{2}x-g_{3}$ with $g_{2},g_{3}\in K$. Let $L$ denote the period lattice corresponding to the fixed Weierstrass form.
We denote the once-punctured elliptic curve associated to $E$ by $X\coloneqq E-O$. We denote the pro-$p$ geometric fundamental group $\pi_{1}(\bar{X},\vec{O})^{(p)}$ with respect to the Weierstrass tangential basepoint at $O$ [Nak99, (2.7) Case 2] by $\Pi_{1,1}$, and identify $\Pi_{1,1}$ with a free pro-$p$ group of rank two with basis $\{\alpha,\beta\}$ in such a way that the commutator $[\alpha,\beta]$ generates an inertia subgroup at $O$.
Note that the pro-$p$ geometric fundamental group $\Pi_{1,0}$ of $E$ is isomorphic to $\Pi_{1,1}^{\rm{ab}}$ through a homomorphism $\Pi_{1,1}^{\rm{ab}}\to\Pi_{1,0}$ induced by the inclusion $X\subset E$. Moreover, there is a natural isomorphism $T_{p}(E)\xrightarrow{\sim}\Pi_{1,0}$. Throughout this paper, we work with the following basis $\{x_{1},x_{2}\}$ of $\Pi_{1,1}$, which is useful when considering the outer Galois action on $\Pi_{1,1}$:
@@
    Let $(\omega_{1,n})_{n\geq 1}$ (resp. $(\omega_{2,n})_{n\geq 1}$) denote the image of $x_{1}$ (resp. $x_{2}$) in $\Pi_{1,1}^{\mathrm{ab}}\cong T_{p}(E)=T_{\mathfrak{p}}(E)\oplus T_{\bar{\mathfrak{p}}}(E)$. Then $(\omega_{1,n})_{n\geq 1}$ (resp. $(\omega_{2,n})_{n\geq 1}$) is a basis of $T_{\mathfrak{p}}(E)$ (resp. $T_{\bar{\mathfrak{p}}}(E)$).
@@
    $z\coloneqq[x_{2},x_{1}]$ generates the inertia subgroup $\langle[\beta,\alpha]\rangle$.
@@
By a result of Kaneko [Kan89, Proposition 2], there exists an automorphism $\gamma$ of $\Pi_{1,1}$ such that the image of $\gamma(\alpha)$ (resp. $\gamma(\beta)$) in $\Pi_{1,1}^{\mathrm{ab}}$ generates $T_{\mathfrak{p}}(E)$ (resp. $T_{\bar{\mathfrak{p}}}(E)$) and the subgroup generated by $\gamma([\beta,\alpha])$ is conjugate to $\langle[\beta,\alpha]\rangle$. Since the inner automorphism group acts trivially on the maximal abelian quotient $\Pi_{1,1}^{\mathrm{ab}}$, we may assume that $\langle\gamma([\beta,\alpha])\rangle=\langle[\beta,\alpha]\rangle$. Now $x_{1}\coloneqq\gamma(\alpha)$ and $x_{2}\coloneqq\gamma(\beta)$ satisfy both conditions. ∎
2.1. Analogue of Ihara’s power series and Soulé characters
In this subsection, we introduce a certain power series which expresses the Galois action on the maximal metabelian quotient of $\Pi_{1,1}$, and Nakamura’s explicit formula of that power series. Then we introduce some non-vanishing results of the coefficients of the power series.
This power series is an elliptic analogue of the universal power series for Jacobi sums introduced by Ihara [Iha86], expressing the Galois action on the metabelian pro-$p$ fundamental group of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$. We will not go into detail about the case of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$ here, and interested readers are encouraged to refer to the pioneering paper of Ihara [Iha86].
First, we define
$\tilde{\Gamma}_{1,1}\coloneqq\{f\in\mathrm{Aut}(\Pi_{1,1})\mid\text{$f$ preserves the conjugacy class of the inertia subgroup $\langle z\rangle$}\},$
its subgroup
$\Gamma_{1,1}^{{\dagger}}\coloneqq\{f\in\mathrm{Aut}(\Pi_{1,1})\mid\text{$f$ preserves $\langle z\rangle$}\},$
and a quotient $\Gamma_{1,1}\coloneqq\tilde{\Gamma}_{1,1}/\mathrm{Inn}(\Pi_{1,1})$ which is called the pro-$p$ mapping class group of type $(1,1)$. The group $\tilde{\Gamma}_{1,1}$ comes equipped with a descending central filtration $\{F^{m}\tilde{\Gamma}_{1,1}\}_{m\geq 1}$ called the weight filtration, which is defined by
@@ 2.1. Analogue of Ihara’s power series and Soulé characters
$F^{m}\tilde{\Gamma}_{1,1}\coloneqq\mathrm{ker}\left(\tilde{\Gamma}_{1,1}\to\mathrm{Aut}(\Pi_{1,1}/\Pi_{1,1}(m+1))\right)$
for every $m\geq 1$. The weight filtration naturally induces filtrations on $\Gamma^{{\dagger}}_{1,1}$ and $\Gamma_{1,1}$ by
$F^{m}\Gamma^{{\dagger}}_{1,1}\coloneqq\Gamma^{{\dagger}}_{1,1}\cap F^{m}\tilde{\Gamma}_{1,1}\quad\text{and}\quad F^{m}\Gamma_{1,1}\coloneqq F^{m}\tilde{\Gamma}_{1,1}\mathrm{Inn}(\Pi_{1,1})/\mathrm{Inn}(\Pi_{1,1}),$
respectively.
@@
Since the intersection $\cap_{m\geq 1}\Pi_{1,1}(m+1)$ of the descending central series is trivial, the intersection $\cap_{m\geq 1}F^{m}\tilde{\Gamma}_{1,1}$ is also trivial. Moreover, the intersection $\cap_{m\geq 1}F^{m}\Gamma_{1,1}$ is also trivial by a result of Asada [Asa95, Theorem 2].
Note that the normalizer subgroup of $\langle z\rangle$ in $\Pi_{1,1}$ is equal to $\langle z\rangle$ itself. It follows that the intersection $F^{m}\Gamma_{1,1}^{{\dagger}}\cap\mathrm{Inn}(\Pi_{1,1})$ is trivial for every $m\geq 3$ and is equal to $\langle\mathrm{inn}(z)\rangle$ for $m=1,2$. Therefore, the natural projection $F^{m}\Gamma_{1,1}^{{\dagger}}\to F^{m}\Gamma_{1,1}$ is an isomorphism for every $m\geq 3$ and, it is surjective with $\langle\mathrm{inn}(z)\rangle$ as its kernel for $m=1,2$ [Nak95, (4.4)].
We define two subgroups $\Psi^{{\dagger}}_{1}\subset\Psi^{{\dagger}}$ of $\mathrm{Aut}\left(\Pi_{1,1}/[\Pi_{1,1}(2),\Pi_{1,1}(2)]\right)$ by

### v3

@@ 1. Introduction
$\chi_{1}\colon G_{K}\to\operatorname{Aut}(T_{\mathfrak{p}}(E))\cong\mathbb{Z}_{p}^{\times}\quad\text{and}\quad\chi_{2}\colon G_{K}\to\operatorname{Aut}(T_{\bar{\mathfrak{p}}}(E))\cong\mathbb{Z}_{p}^{\times}.$
@@
Note that $\chi_{1}\chi_{2}$ coincides with the $p$-adic cyclotomic character $\chi_{\mathrm{cyc}}$. Let $\boldsymbol{m}=(m_{1},m_{2})$ be a pair of integers. We define the character $\chi^{\boldsymbol{m}}$ by
$\chi^{\boldsymbol{m}}\coloneqq\chi_{1}^{m_{1}}\chi_{2}^{m_{2}}\colon G_{K}\to\mathbb{Z}_{p}^{\times}.$
Note that it factors through $\operatorname{Gal}(K(p^{\infty})/K)$ if $m_{1}\equiv m_{2}\bmod\lvert O_{K}^{\times}\rvert$.
Let $M$ be a $\mathbb{Z}_{p}$-module on which $G_{K}$ acts continuously. we denote the $\chi^{\boldsymbol{m}}$-twist of $M$ by $M(\boldsymbol{m})$. Note that the $\mathfrak{p}$-adic (resp. $\bar{\mathfrak{p}}$-adic) Tate module $T_{\mathfrak{p}}(E)$ is isomorphic to $\mathbb{Z}_{p}(1,0)$ (resp. $\mathbb{Z}_{p}(0,1)$), and $\mathbb{Z}_{p}(m,m)$ is simply the $m$-th Tate twist $\mathbb{Z}_{p}(m)$.
@@ 2. Preliminaries
In this section, we prepare backgrounds which are necessary to state the main result of this paper (Theorem 2.14). We explain:
  • a certain basis $\{x_{1},x_{2}\}$ of the pro-$p$ geometric fundamental group of once-punctured elliptic curve $X$ (Lemma 2.1),
  • an analogue of the Deligne-Ihara conjecture (Conjecture 2.3),
  • an elliptic analogue of the universal power series for Jacobi sums [Iha86], the explicit formula for that power series [Nak95], conditional nonvanishing of certain Kummer characters associated to the power series [Ish25] and a reformulation of Conjecture 2.3 in terms of these characters (Proposition 2.11),
  • the main result (Theorem 2.14) in a precise manner.
In the rest of the present paper, let $K$ be an imaginary quadratic field of class number one, and $p\geq 5$ a prime which splits in $O_{K}$ as $(p)=\mathfrak{p}\bar{\mathfrak{p}}$. Moreover, let $E$ be an elliptic curve defined over $K$ having complex multiplication by $O_{K}$ with a fixed Weierstrass form of $y^{2}=4x^{3}-g_{2}x-g_{3}$ with $g_{2},g_{3}\in K$. Let $L$ denote the period lattice corresponding to this form.
We denote the once-punctured elliptic curve associated to $E$ by $X\coloneqq E-O$, where $O$ is the origin of $E$. We denote the pro-$p$ geometric fundamental group $\pi_{1}(\bar{X},\vec{O})^{(p)}$ with respect to the Weierstrass tangential basepoint at $O$ [Nak99, (2.7) Case 2] by $\Pi_{1,1}$. Note that $\Pi_{1,1}$ is a free pro-$p$ group of rank two, and we have an injective homomorphism $\mathbb{Z}_{p}(1)\subset\Pi_{1,1}$ associated with the tangential basepoint $\vec{O}$. Moreover, the pro-$p$ geometric fundamental group of $E$ is isomorphic to $\Pi_{1,1}^{\rm{ab}}$ through a natural homomorphism induced by the inclusion $X\subset E$, and there is a natural isomorphism $T_{p}(E)\xrightarrow{\sim}\Pi_{1,0}$. Throughout this paper, we work with the following basis $\{x_{1},x_{2}\}$ of $\Pi_{1,1}$, which is useful when considering the outer Galois action on $\Pi_{1,1}$:
@@
    Let $(\omega_{1,n})_{n\geq 1}$ (resp. $(\omega_{2,n})_{n\geq 1}$) denote the image of $x_{1}$ (resp. $x_{2}$) in $\Pi_{1,1}^{\mathrm{ab}}\cong T_{p}(E)=T_{\mathfrak{p}}(E)\oplus T_{\bar{\mathfrak{p}}}(E)$. Then $(\omega_{1,n})_{n\geq 1}$ (resp. $(\omega_{2,n})_{n\geq 1}$) generates $T_{\mathfrak{p}}(E)$ (resp. $T_{\bar{\mathfrak{p}}}(E)$).
@@
    $z\coloneqq[x_{2},x_{1}]$ generates the inertia subgroup $\mathbb{Z}_{p}(1)\subset\Pi_{1,1}$ determined by $\vec{O}$.
@@
By a result of Kaneko [Kan89, Proposition 2], we can choose a basis $\{x_{1},x_{2}\}$ so that it satisfies (1) and the subgroup generated by $x_{1},x_{2}]$ is conjugate to the inertia subgroup determined by $\vec{O}$. Since the inner automorphism group acts trivially on the maximal abelian quotient $\Pi_{1,1}^{\mathrm{ab}}$, we may replace $\{x_{1},x_{2}\}$ with its conjugate so that it also satisfies (2). This concludes the proof. ∎
2.1. Structure of graded Lie algebra
In this subsection, we introduce a certain filtration on the absolute Galois group $G_{K}$ and formulate a working hypothesis on the structure of the graded Lie algebra associated to this filtration (Conjecture 2.3). This is an analogue of the so-called Deligne-Ihara conjecture in the case of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$. We then study some fundamental properties of the fixed field of $\operatorname{Ker}(\rho_{X,p})$ of the pro-$p$ outer Galois representation associated to $X$, and state the main result.
First, we introduce the pro-$p$ mapping class group of type $(1,1)$: Denote
[…]
and define the pro-$p$ mapping class group of type $(1,1)$ by the quotient
$\Gamma_{1,1}\coloneqq\tilde{\Gamma}_{1,1}/\operatorname{Inn}(\Pi_{1,1})\subset\operatorname{Out}(\Pi_{1,1}).$
We define the weight filtration on $\tilde{\Gamma}_{1,1}$ and $\Gamma_{1,1}$ by
@@ 2.1. Structure of graded Lie algebra
$F^{m}\tilde{\Gamma}_{1,1}\coloneqq\operatorname{Ker}\left(\tilde{\Gamma}_{1,1}\to\operatorname{Aut}(\Pi_{1,1}/\Pi_{1,1}(m+1))\right)\quad\text{and}\quad F^{m}\Gamma_{1,1}\coloneqq F^{m}\tilde{\Gamma}_{1,1}\operatorname{Inn}(\Pi_{1,1})/\operatorname{Inn}(\Pi_{1,1})$
@@
for each positive integer $m$. They define descending central filtrations, and the intersection $\cap_{m\geq 1}F^{m}\tilde{\Gamma}_{1,1}$ is trivial since $\cap_{m\geq 1}\Pi_{1,1}(m+1)$ is trivial. Moreover, the intersection $\cap_{m\geq 1}F^{m}\Gamma_{1,1}$ is also trivial by a result of Asada [Asa95, Theorem 2].
The absolute Galois group $G_{K}$ inherits a descending central filtration $\{F^{m}G_{K}\}_{m\geq 1}$ by taking a pullback $F^{m}G_{K}\coloneqq\rho_{X,p}^{-1}(F^{m}\Gamma_{1,1})$, and we call it the weight filtration (on $G_{K}$). Note that we have $F^{1}G_{K}=G_{K(E[p^{\infty}])}$. We form graded quotients and their directed sum as
Since the weight filtration is descending and central, the direct sum $\mathfrak{g}$ is a graded Lie algebra over $\mathbb{Z}_{p}$ whose bracket is induced by commutators.
We briefly summarize properties of the graded quotients:
Proposition 2.2.
Let $m$ be a positive integer. The following assertions hold:
  1. (1)
    $\mathfrak{g}_{m}$ is trivial whenever $m$ is odd, and it is also trivial when $m=2$.
  2. (2)
    $\mathfrak{g}_{m}$ is a free $\mathbb{Z}_{p}$-module of finite rank.
  3. (3)
    $\mathfrak{g}_{m}\otimes\mathbb{Q}_{p}$ is isomorphic to a direct sum of (a direct sum of) $\mathbb{Q}_{p}(\boldsymbol{m})$’s as $G_{K}$-modules, where $\boldsymbol{m}$ runs over pairs of positive integers satisfying $|\boldsymbol{m}|=m$. In particular, the graded Lie algebra $\mathfrak{g}\otimes\mathbb{Q}_{p}$ has a structure of a bigraded Lie algebra.
Proof.
(1) The assertion follows from [Nak95, (4.2) Proposition] and [Nak95, (4.4)]. (2) The assertions follows since $\mathfrak{g}_{m}$ is embedded into $F^{m}\Gamma_{1,1}/F^{m+1}\Gamma_{1,1}$, which is a free $\mathbb{Z}_{p}$-module [NT93, Corollary (1.16), $(\mathrm{ii})$] of finite rank. (3) The assertion follows immediately from the $\operatorname{GL}_{2}(\mathbb{Z}_{p})$-equivariance of the commutative diagram given in [NT93, Theorem (1.14)]. ∎
Remark 2.2.1.
The Lie algebra $\mathfrak{g}\otimes\mathbb{Q}_{p}$ only depends on the isomorphism class of $\bar{X}$ not on $X$: There is only one isomorphism class of elliptic curves over $\bar{K}$ having complex multiplication by $O_{K}$ and a base change by a finite extension preserves $\mathfrak{g}\otimes\mathbb{Q}_{p}$.
The Deligne-Ihara conjecture states that, a $\mathbb{Q}_{p}$-graded Lie algebra, which is defined similarly to $\mathfrak{g}\otimes\mathbb{Q}_{p}$ using the pro-$p$ outer Galois representation of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$ instead that of $X$, is freely generated by one element in each odd degree $m\geq 3$. This conjecture appears in e.g. [Iha02, p.251, after Theorem (1)] and [Sha02, the discussion before Theorem 1.1], and there the conjecture was attributed to Deligne’s monograph [Del89]. We formulate an analogue of this conjecture, as follows. Let us define the set of pairs of integers
$I\coloneqq\{\boldsymbol{m}=(m_{1},m_{2})\in\mathbb{Z}_{\geq 1}^{2}\setminus\{\boldsymbol{1}\}\mid m_{1}\equiv m_{2}\bmod\lvert O_{K}^{\times}\rvert\}$.
This set plays a similar role as the set $1+2\mathbb{Z}_{\geq 1}=\{3,5,7,\dots\}$ appearing in the case of the projective line minus three points. Our working conjectural description of the Lie algebra $\mathfrak{g}\otimes\mathbb{Q}_{p}$ is as follows:
Conjecture 2.3.
$\mathfrak{g}\otimes\mathbb{Q}_{p}$ is freely generated by one element in each degree in $I$.
Remark 2.3.1.
(1) One can prove the generation portion of Conjecture 2.3, assuming that the second cohomology group $H^{2}_{\mathrm{\acute{e}t}}(O_{K}[1/p],\mathbb{Z}_{p}(\boldsymbol{m}))$ is finite for every $\boldsymbol{m}\in I$, by using (a slightly modified version of) the theory of weighted completions developed by Hain-Matsumoto [HM03]. This provides one evidence for Conjecture 2.3, at least for its generation portion. At the writing of the present paper, however, we do not know how to prove the freeness.
(2) Our proof of Theorem A (given in Section 4) shows that, under the assumption of Theorem A, one can even choose a basis of $\mathfrak{g}\otimes\mathbb{Q}_{p}$ so that it also freely generates $\mathfrak{g}$, cf. Corollary 4.14 (1). This integral version of Conjecture 2.3 is expected not to hold in general, and we plan to investigate the integral structure of $\mathfrak{g}$ in future research.
Although the statement of Conjecture 2.3 may seem unexpected at first sight, we rephrase it in a different manner in the next subsection by using certain natural characters arising from the $G_{K}$-action on the metabelian fundamental group. The author thinks it gives a more natural statement.
2.2. Elliptic analogues of Soulé characters
The aim of this subsection is twofold: First, we study the $G_{K}$-action on the maximal metabelian quotient of $\Pi_{1,1}$ and present some previous results. Then we interpret Conjecture 2.3 in terms of elliptic analogues of the Soulé characters, which is useful in proving our main result. We introduce a certain power series that expresses the Galois action on the maximal metabelian quotient of $\Pi_{1,1}$: This power series is an elliptic analogue of the universal power series for Jacobi sums introduced by Ihara [Iha86], expressing the Galois action on the metabelian pro-$p$ fundamental group of $\mathbb{P}^{1}_{\mathbb{Q}}-\{0,1,\infty\}$. It is extensively studied in the paper of Nakamura [Nak95], and there an application to anabelian geometry of once-punctured elliptic curves is given. We will not go into detail here, and interested readers are encouraged to refer to the papers [Iha86] (in the case of genus zero) [Nak95] (in the case of genus one).
First, we define the subgroup of $\tilde{\Gamma}_{1,1}$ by